OpenStudy (anonymous):

Find the max and min values of f (x, y) = cos t^2 + sin t^2 − cos t− sin t

OpenStudy (anonymous):

F(x,y)=1-cost-sint but i do not know where to go from here

OpenStudy (amistre64):

f(x,y) = an equation with no xs or ys ???

OpenStudy (amistre64):

cos^2 + sin^2 = 1 regardless

OpenStudy (amistre64):

i read that wrong tho lol

OpenStudy (anonymous):

the original equation is : f (x, y) = x2 + y2 − x − y on the closed unit disk D : x^2+y^2 ≤ 1. (Hint: Recall that the unit circle x^2+y^2 = 1 can be parametrized as x = cos t, y = sin t.) so i just substituted x and y values

OpenStudy (anonymous):

i have to find the max and min values

OpenStudy (amistre64):

you need to find your partials right?

OpenStudy (anonymous):

yes and i have done that i have found the local min value (1/2,1/2)

OpenStudy (anonymous):

i just dont know how to find the max

OpenStudy (amistre64):

the links aint up and running so youll have to copy paste this to see http://www.wolframalpha.com/input/?i=+f+%28x%2C+y%29+%3D+x2+%2B+y2+%E2%88%92+x+%E2%88%92+y

OpenStudy (amistre64):

it has no max; just like a parabola on an x,y plane has no max

OpenStudy (amistre64):

f(x, y) = x2 + y2 −x−y df/dx = 2x −1 df/dy = 2y -1 when 2x-1=0 and 2y-1=0, you have "normal" vector that point up to indicate a min or max of the surface

OpenStudy (amistre64):

or at least a critical point, it could also be a saddle back but not this time